### v1

# FTPI-MINN-24-02, UMN-TH-4310/44

Abstract
The first Edition of this book was released in 2000, just before the symposium “Thirty Years of Supersymmetry” was held at the William I. Fine Theoretical Physics Institute (FTPI) of the University of Minnesota. Founders and trailblazers of supersymmetry descended on FTPI, as well as a large crowd of younger theorists deeply involved in research in this area. Since then 23 years have elapsed and significant changes happened in supersymmetry (SUSY). Its history definitely needs an update. Such an update is presented below. The Second Edition of the revised collection will be released in 2024 by WS.
The first Edition of this book was released in 2000, just before the symposium “Thirty Years of Supersymmetry” was held at the William I. Fine Theoretical Physics Institute (FTPI) of the University of Minnesota. Founders and trailblazers of supersymmetry descended on FTPI, as well as a large crowd of younger theorists deeply involved in research in this area. Remarkably, it was at this event that many of the early pioneers of the field met face-to-face for the first time. Table 1 presents the Contents of the historical part of the SUSY-30 Proceedings [2]. Since then 23 years have elapsed and significant changes happened in supersymmetry (SUSY). Its history definitely needs an update.
Below the reader will see a table which can be viewed as a starting point for the current edition. In 2000 supersymmetry explorations were on the rise and SUSY-based phenomenology attracted hundreds of young researchers, with around 2600 original papers published annually (see the graphs 3, 4 on pages 14, 15). Expectations of an imminent experimental discovery of superpartners were prevailing in the community; the general mood was rosy.
Table 1: The historic part in [2]. Table of contents
Part 1: Supersymmetry from the East N. Koretz-Golfand Supersymmetry – 30 E.P. Likhtman Around SUSY 1970 V. Akulov Nonlinear way SUSY and ${\cal N}$ extended SUSY S.I. Volkova, A.A. Zheltukhin Glimpses of Dmitry Volkov life and work V.A. Soroka Supersymmetry and the odd Poisson bracket Part 2: Supersymmetry from the West Pierre Ramond Boson-fermion confusion: The string path to SUSY J.H. Schwarz String theory origins of supersymmetry J.-L. Gervais Symmetries: early days and nowadays B. Sakita Symmetries of fermions in the lowest Landau level P. Fayet About the origins of supersymmetric standard model J. Iliopoulos Non-renormalization theorems in globalsupersymmetry L. O’Raifeartaigh Chiral fermions on the lattice P. West Non-renormalization theorems in supersymmetric theories M.F. Sonius Recollections of a young contributor S. Ferrara Superconformal algebras and supergravities in higher dimension A.H. Chamseddine, R. Arnowitt, P. Nath Supergravity unification Bruce de Witt Elrctric-magnetic dualities in supergravity
Figure 1: The superworld tree in 2023
The modern perspective on supersymmetry looks different. So far, all searches of SUSY in LHC experiments have brought negative results, a great disappointment. The rosy mood of the early 2010s faded away.
On the other hand, the unique power of supersymmetry became obvious in mathematical physics, string mathematics, and such topics as black holes and critical (BPS) solitons, especially at strong coupling where SUSY hardly has any competitors. For the first time ever, the analytic proof of quark confinement was found in ${\cal N}=2$ super-Yang-Mills theory [3] slightly deformed by an ${\cal N}=1$ perturbation.
This is reflected in Fig. 1 which I designed especially for the current edition. (See also the graph in Fig. 5 on page 16). Figure 1 demonstrates that the “Pheno” branch of supersymmetry is drying out – its future is unclear at the moment. The “strong coupling” branch dealing with “imaginary” worlds and more mathematical applications of SUSY continues to grow.
“Supersymmetry is so beautiful and suggestive that most of us think it has got to show up sometime in nature, although so far it hasn’t,” said Steven Weinberg in his 1979 Nobel speech. I still believe in Weinbeg’s prophecy but it is reasonable to ask “when?” Unfortunately, I am not a prophet and the only answer I can give is “The future’s not ours to see…”
“We cannot guess the way of word
In real world, how it’ll return…”  [^2] A quotation from a Russian poet Fyodor Tyutchev (1803-1873). Translated by Lyudmila Purgina.
In this volume we collected personal memories describing days long gone, with their excitement and even euphoria. These testimonies were penned by the discovers of supersymmetery, trailblazers, and pioneers themselves. Hopefully they will be used by future scholars, in particular, historians of science. Unfortunately, even in 2000 when the First Edition was being prepared it was too late to obtain personal memoirs from Yuri Golfand (1922-1994) and Abdus Salam (1926-1996). We had to settle for their collaborators – Evgeny Likhtman in the first case and John Strathdee in the second. [^3] Here I have to make a few explanatory remarks. (i) In English literature Golfand’s surname is written in two different ways – Golfand or Gol’fand. The apostrophe in the latter version indicates that a certain “mute” letter of the Russian alphabet is omitted. Phonetically this omitted letter makes the previous “l” soft, as in French; (ii) Evgeny Likhtman was Golfand’s PhD student and a co-author of three crucial papers; (iii) John Strathdee was Salam’s co-author of 27 years! He retired from ICTP in the late 1990s and hid somehwhere in New Zealand. Finding him there was not easy; I was told that he had left his mailing address only to a certain secretary at ICTP. I contacted her and eventually she entrusted me with Strathdee’s address after I promised her to use it just once. Strathdee’s reply to my letter contained seven lines, for which I am grateful, see Chapter 5, page 246.
Almost all of the early explorers of supersymmetry, its founding fathers, have left our world after 2000. This is a sad – but alas…– a natural process. Julius Wess and Bruno Zumino, who (among others) were the masterminds and discoverers of four-dimensional supersymmetry, died in 2007 and 2014 respectively. They were lifelong friends and collaborators. I feel that Julius and Bruno, who gave so much to our community, deserve a special “In Memoriam” chapter in the current edition (see Chapter  1).
The available literature about the dawn of supersymmetry has significantly expanded since 2000. A wealth of information is presented in [2] and the accompanying volume [4]. In 2004, a fascinating “Concise Encyclopedia of Supersymmetry” was published [5] with more than 700 articles and a historical section written specifically for this Encyclopedia by E. Likhtman, Dmitry Volkov, V. Akulov, H. Miyazawa, G. Stavraki, V. Kac, V. Pakhomov, J. Łopuszanski, R. Haag, and D. Leites. Next, it is worth mentioning the collection [6] narrating the story of Felix Berezin, an outstanding mathematician who created a mathematical apparatus used in supersymmetric field theories. The integral over the anticommuting Grassmann variables that he introduced in the 1960s paved the way for the path integral formulation of quantum field theory with fermions, the heart of modern supersymmetric field theories and superstrings. [^4] See Marinov’s article in Chapter 6, page 302. The Berezin integral is named for him, as is the closely related construction of the Berezinian, which may be regarded as the superanalogue of the determinant. In the same year the Dmitry Volkov Memorial Volume [7] was published in Ukraine in Russian. It contains both recollections of his colleagues and friends and Volkov’s selected interviews spanning the years 1986-1991. Two detailed essays on Yuri Golfand appeared in Part 2 of the book [8].
In 2012 a profound collection The Birth of String Theory [9] was released. It contains a wealth of information on the early days of string theory some of which is relevant to our Chapter 2; see in particular Ramond’s articles.
Approximately a year before his death Stanley Deser published a book of recollections about his life and work [10]. The chapter entitled “A Big Year” will be especially interesting to the reader since it narrates Deser’s viewpoint on the story of the creation of Supergravity (SUGRA). Other creators of SUGRA published a number of reviews containing their recollections. The appropriate excerpts will be presented in Chapter 4 of the current Edition.
Finally, a remarkable connection between the late works of W. Heisenberg and the discovery of nonlinearly realized supersymmetry by Volkov and Akulov was revealed in [11].
Compared to the 2000 edition, the Second Edition is significantly expanded. First, as was mentioned, we add Chapter 1, In Memoriam: Julius Wess and Bruno Zumino, which presents recent recollections on Zumino and Wess. Chapter 4, Local Supersymmetry (Supergravity), is completely revised, and now includes a number of articles which appeared after 2000. Chapter 7 containing a brief report on SUSY-50  [^5] The conference 50 Years of Supersymmetry (SUSY 50) was organized by FTPI, University of Minnesota, on May 18-20, 2023. was prepared specifically for the Second Edition. In the newly organized Appendix the reader can find English translations of some of the early papers by Golfand, Likhtman, Volkov, Akulov, and Soroka which were published in JETP and therefore are not easily accessible. In a number of instances I added comments or explanatory remarks, see e.g. page 429.
Other additions include an essay of John Iliopoulos on the early results, dreams and expectations of supersymmetry, as well as a brief review of the present situation (Chapter 5). In the same Chapter the reader will find recollections about Victor Ogievetsky (1928-1996), the creator of the harmonic superspace for ${\cal N}=2$ theories. Upon reflection, I decided that supersymmetry at strong coupling should be viewed as a “special branch” of SUSY (Fig. 1), and a few words about its pioneers are in order in the Second Edition. Correspondingly, I added an interview with Nati Seiberg in Chapter 5 devoted to the pioneers, as well as my commentary.
[…]
I also correct two mistakes. The Foreword to the First Edition stated:
Often students ask where the name “supersymmetry” came from. It seems that it was coined in the paper by Salam and Strathdee [12] where these authors constructed supersymmetric Yang–Mills theory [using superspace formalism]. This paper was received by the editorial office on June 6, 1974, exactly eight months after that of Wess and Zumino. Super-symmetry (with a hyphen) is in the title, while in the body of the paper Salam and Strathdee use both, the old version of Wess and Zumino, “super-gauge symmetry,” and the new one.
When Bruno Zumino read this paragraph (it was long ago) he got upset and replied me in a letter which also contained a typewritten text of his talk [13].
He wrote:
Dear Misha,
I did the same in my talk at 17th International Conference on High-Energy Physics – seemingly – a few days before Salam and Strathdee. Please, read it.
I apologized and promised Bruno to correct my mistake. Now, many years later I fulfill the promise. The reader can find Bruno’s 1974 talk added in Chapter 3, just after Julius Wess’ talk.
Another unintentional omission was pointed out to me by Pierre Ramond. He suggested supplementing the narrative on Pierre Fayet’s important contribution which eventually led to MSSM in the First Edition,
In the subsequent years, Pierre Fayet combined the Brout-Englert-Higgs mechanism with supersymmetry and introduced superpartners to the Standard Model particles [14]. This was the inception and birth of the minimal supersymmetric standard model (MSSM). Heavy quarks unknown in the 1970s neutrino mixing angles and some details were introduced somewhat later […],
by the reference to Dimopoulos and Georgi [15] who developed a supersymmetric model with softly broken SUSY in the context of Grand Unification in 1981.
Figure 2: One of the posters honoring Bruno Zumino’s 90th anniversary.
Preparing the Second Edition I noticed a change in the community which is social rather than scientific. In the 1970, 1980s, and 1990s only a few early researchers in this area were aware of the contributions of Golfand, Likhtman, Volkov, Akulov and Soroka. Virtually none of the young theorists knew about works on four-dimensional supersymmetry and supergravity “before Wess-Zumino.” This was one of the dark consequences of the 70-year Iron Curtain between the USSR and the rest of the world erected by the communist regime. The proceedings of SUSY-30 [2], the subsequent publications (the corresponding references can be found on page References) and a number of talks on the origins of SUSY I gave in the last two decades on various occasions led to a general recognition of the “contribution from the East.” Now it has become common knowledge  [^6] See e.g. Chapter 2, Ramond’s essay on page 78. which I view as my modest achievement in restoring “fair attribution” practice in our community.
Concluding this Foreword I have to mention that, unfortunately, Gordy Kane was unable to participate in preparation of the current edition. Therefore, I am solely responsible for all inaccuracies the reader may find in the added material.
Acknowledgments
With great pleasure I acknowledge many illuminating discussions with Dan Freedman and Pierre Ramond.
I am very grateful to Lakshmi Narayanan – my World Scientific editor for many years – with whom I worked on around a dozen projects. Now, when she is retiring I would like to say thank you, Lakshmi.
This work is supported in part by DOE grant DE-SC0011842.